Karlottalotta:

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20

SaltTheLoser:

\[(x^2+2x)+(y^2+4y) = 20\\(2/2)^2 = 1\\(4/2)^2=4\\(x^2+2x+1)+(y^2+4x+4)=20+1+4\\(x+1)^2+(y+2)^2=5^2\\(x-h)^2+(y-k)^2=r^2\\(-1,-2)~~~r=5\]